### v2

@@ 4.4 ROI Classification
For patch-level tissue classification tasks, we evaluate the transfer performance and representation ability of different foundation models using a linear probe, inspired by the approach employed in DINOv2. [56, 84]. Initially, we extract features from the images using the pretrained foundation models. Subsequently, we employ a linear layer for performing classification. To optimize the model, we utilize AdamW [85] with an initial learning rate of 5e-4 and weight decay of 1e-5. Additionally, we incorporate a cosine annealing scheduler to update the learning rate during training [86]. In order to obtain the best model, we set the maximum number of epochs to 3000 and implemented early stopping with patience of 100 epochs. For ensuring fair comparison, we maintain a consistent batch size of 256 across all methods.
To evaluate the performance of patch-level tissue classification, we consider the impact of class imbalance in the dataset and assess the metrics of balanced accuracy, weighted F1 score, and AUC. These metrics provide comprehensive insights into the classification performance, accounting for both accuracy and the ability to handle imbalanced class distributions. Specifically, we compare the foundation models across 12 tasks, including (1) Colorectal cancer (CRC) tissue classification, (2) Clear cell colorectal cancer (CCRCC) tissue classification, (3) Breast cancer tissue classification, (4) CRC polyp classification, (5) Microsatellite instability screening, (6) PanCancer tissue classification, (7) Tumor infiltrating lymphocyte classification, (8) Esophageal carcinoma subtyping, (9) Metastatic tissue classification, (10) Lung adenocarcinoma tissue classification, (11) Breast cancer image classification, and (12) Colon tissue classification. For all experiments in this section, we estimate the model performance using non-parametric bootstrapping with 1,000 bootstrap replicates. We employ Torchmetrics [87] for bootstrapping sampling and obtain the mean and standard deviation of the metrics. The experimental results are presented in Extended Data Table 21 to Extended Data Table 32. Furthermore, we report the average performance of the patch-level tissue classification results across 12 tasks in Table 20, demonstrating the superior performance of GPFM.
@@
PanCancer-TIL for TIL classification (2 classes). This dataset [93, 94] is collected for tumor infiltrating lymphocyte (TIL) classification including 304,097 images with a size of 100100 pixels at 0.5 micrometers per pixel.
@@
The experimental results are presented in Extended Data Table 31.
@@
The experimental results are presented in Table 32.
4.5 Pathological Tissue Retrieval.
In the linear probe evaluation tasks, we extract semantically-rich features using different foundation models and then construct a task-specific classifier. These features are not only applicable for supervised learning but also prove to be valuable for image-to-image retrieval. The primary goal of this application is to retrieve images that share the same class label as a given query image, thereby facilitating efficient image retrieval. To initiate the pathological tissue image retrieval process, we begin by embedding all images using pretrained foundation models. Next, each image in the test set is treated as a query and compared against the images in the training set. To ensure that all features have a comparable impact on the computation of similarity, we independently normalize each feature component to the range [0, 1] [101]. This normalization process involves calculating the mean and variance of the training set features, which are then used to normalize both the training and testing features.
@@ 4.6 Pathology Visual Question Answering
The objective of this subsection is to evaluate the performance of our proposed pathology foundation model in the context of Visual Question Answering (VQA) tasks. To achieve this, we utilize the PathVQA dataset [62], a widely recognized and extensively used benchmark in the pathology domain, as the evaluation dataset for our experiments. This dataset provides a comprehensive assessment of the model’s ability to comprehend and reason about visual pathology information, enabling accurate answering of queries related to observed features. To ensure a rigorous and comparative analysis, we adopt the Multi-modal Unified Medical Captioning (MUMC) method [102], which currently represents the state-of-the-art approach on the PathVQA dataset. The MUMC method has exhibited superior performance in leveraging the synergies between visual and textual information for medical image understanding tasks.
The VQA model architecture consists of four main components: the image encoder, text encoder, multimodal encoder, and answering decoder. The image encoder is responsible for capturing domain-specific visual features. We employ various pathology foundation models as the image encoder. During the fine-tuning process, the weights of the image encoder are kept frozen to preserve the integrity of the pre-trained visual representations and focus on learning task-specific multimodal interactions.
@@ 4.7 Pathology Report Generation
The task of pathology report generation is inspired by existing works on Chest X-ray and other medical report generation [104, 105, 106]. In this task, the report generation model takes a WSI as input and generates the corresponding pathology report. Specifically, the input WSI is first processed by foundation models to extract an initial representation. This representation is then fed into the encoder-decoder architecture of report generation models to produce the decoded pathology report. During this process, the visual encoder further processes the initial representations of WSIs through specific designs [105, 106, 63] to obtain the optimal WSI features for the report decoding stage. The text decoder of the model then utilizes these features for report generation. A good initial representation of WSI could significantly facilitate both the visual encoding and textual decoding stages. Consequently, the quality of the generated report is directly influenced by the representations provided by the foundation models. In this task, we adopt the HistGen model [63] for WSI report generation and set the learning rate to 1e-4, and weight decay to 0.8 per epoch. The model is trained for 40 epochs with batch size 1 using features extracted from different foundation models.
To evaluate the report generation performance of foundation models, we utilize natural language generation metrics including BLEU [107], METEOR [108], and ROUGE-L [109], in which BLEU is further split into BLEU-1, BLEU-2, BLEU-3, and BLEU-4 for evaluation of different granularity. These metrics provide a robust framework for evaluating machine-generated text, each bringing unique strengths to assess different aspects of text quality. This task is conducted on the TCGA WSI-Report dataset proposed in [63], containing 7,690 WSIs and the paired diagnosis reports in total. A 7:1:2 train-validation-test split is employed and the experimental results are reported in Extended Data Table 35.
@@ 4.8 Computing Software and Hardware
In this project, we utilized PyTorch [110] (version 2.1.2 with CUDA 12.1) for both pretraining and evaluating downstream tasks. To pretrain the GPFM model, we incorporated established foundation models, namely UNI, Phikon, and CONCH, as additional teachers. It is worth noting that access to UNI and CONCH requires a prior application submission. The GPFM model was pretrained using the FullyShardedDataParallel (FSDP) technique on 28 80GB NVIDIA H800 GPU nodes. All other data processing and evaluation for downstream tasks were carried out on a server equipped with 8 NVIDIA RTX 3090 GPUs. To assess the model’s performance, we employed Torchmetrics [87] and Scikit-learn [111] for metric evaluation. For WSI processing, we relied on openslide-python (version 1.2.0) [70] and the CLAM [71] codebase. Pathology VQA evaluation was conducted using the MUMC [102] codebase. Furthermore, for histology report generation, we utilized the HistGen [63] codebase. Please refer to Extended Data Table 42 for a comprehensive list of the aforementioned models and libraries utilized in this study.
@@ 5 Data availability
This study incorporates a total of 47 datasets. Out of these, 33 datasets are utilized for pretraining, and a subset of them is also employed for evaluation purposes. The remaining 14 datasets are specifically dedicated to downstream task evaluation. For detailed information about the data used in this project, please refer to Extended Data Table 43. The splits of the dataset can be found in our GitHub repository.
@@ 7 Ethics declarations
This project has been reviewed and approved by the Human and Artefacts Research Ethics Committee (HAREC). The protocol number is HREP-2024-0212.
@@ 8 Author contributions
J.M. conceived the study and designed the experiments. J.M., Z.G., and F.Z. collected the data for self-supervised learning and downstream task evaluation. J.M. performed foundation model pretraining and conducted patch-level tissue classification tasks. Y.L. and X.J. participated in discussions regarding the design of the self-supervised learning framework and were responsible for reproducing the foundation models. J.M., F.Z., and Y.C. evaluated the weakly-supervised WSI classification task. Z.G. performed survival analysis and report generation tasks. Y.W. and Y.X. conducted pathological image retrieval and curated the data of WSI-report pairs. Z.Z. performed pathological image visual question answering. C.J. assisted in result analysis and the creation of visualized attention maps. J.M. and Z.G. prepared the manuscript with input from all co-authors. R.C.K.C, A.H., and L.L. provided medical guidance. J.W. offered insightful suggestions for the experimental design and thoughtfully directing the research trajectory. K.T.C reviewed and refined the draft. H.C. supervised the research.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62202403), Hong Kong Innovation and Technology Fund (No. PRP/034/22FX), Research Grants Council of the Hong Kong Special Administrative Region, China (No. R6003-22) and HKUST Frontier Technology Research for Joint Institutes with Industry (No. OKT24EG01). We also thank the support of HKUST SuperPod for providing the GPU platform for foundation model pretraining.
@@ Appendix A Extended Data
| Dataset | ResNet50 | Phikon | Ctran. | UNI | CONCH | PLIP | GPFM |
| BRACS-3 | 0.835 | 0.898 | 0.908 | 0.913 | 0.923 | 0.866 | 0.936 |
| Camelyon | 0.922 | 0.967 | 0.957 | 0.987 | 0.965 | 0.929 | 0.988 |
| BRACS-7 | 0.772 | 0.818 | 0.857 | 0.855 | 0.841 | 0.814 | 0.871 |
| BRCA | 0.846 | 0.936 | 0.931 | 0.946 | 0.944 | 0.893 | 0.950 |
| UBC-OCEAN | 0.919 | 0.966 | 0.960 | 0.969 | 0.957 | 0.946 | 0.972 |
| RCC | 0.963 | 0.995 | 0.981 | 0.982 | 0.988 | 0.980 | 0.996 |
| NSCLC | 0.929 | 0.982 | 0.963 | 0.977 | 0.986 | 0.942 | 0.986 |
| CPTAC-NSCLC | 0.937 | 0.967 | 0.965 | 0.960 | 0.961 | 0.939 | 0.974 |
| PANDA | 0.884 | 0.943 | 0.925 | 0.944 | 0.921 | 0.903 | 0.948 |
| LUAD TP53 | 0.742 | 0.841 | 0.770 | 0.766 | 0.836 | 0.821 | 0.855 |
| Glioma IDH-1 | 0.945 | 0.990 | 0.974 | 0.996 | 0.995 | 0.974 | 0.998 |
| EBRAINS | 0.959 | 0.990 | 0.985 | 0.996 | 0.992 | 0.977 | 0.996 |
| BRCA | 0.678 | 0.770 | 0.704 | 0.722 | 0.677 | 0.691 | 0.757 |
| BLCA | 0.696 | 0.692 | 0.670 | 0.690 | 0.644 | 0.665 | 0.711 |
| KIRC | 0.634 | 0.708 | 0.671 | 0.674 | 0.659 | 0.652 | 0.697 |
| KIRP | 0.630 | 0.680 | 0.649 | 0.659 | 0.623 | 0.674 | 0.688 |
| LUAD | 0.646 | 0.738 | 0.633 | 0.673 | 0.641 | 0.628 | 0.739 |
| STAD | 0.636 | 0.705 | 0.669 | 0.688 | 0.636 | 0.657 | 0.694 |
| LUSC | 0.667 | 0.779 | 0.703 | 0.746 | 0.661 | 0.686 | 0.777 |
| COADREAD | 0.707 | 0.680 | 0.691 | 0.679 | 0.653 | 0.708 | |
| GBMLGG | 0.724 | 0.760 | 0.745 | 0.778 | 0.735 | 0.738 | 0.785 |
| LUAD-CPTAC | 0.517 | 0.577 | 0.624 | 0.575 | 0.649 | 0.608 | 0.626 |
| CESC | 0.611 | 0.747 | 0.702 | 0.735 | 0.677 | 0.694 | 0.754 |
| SKCM | 0.699 | 0.773 | 0.743 | 0.747 | 0.721 | 0.709 | 0.777 |
| CRC-100K | 0.983 | 0.993 | 0.995 | 0.991 | 0.993 | 0.990 | 0.995 |
| CRC-MSI | 0.706 | 0.772 | 0.802 | 0.797 | 0.810 | 0.691 | 0.812 |
| PCAM | 0.926 | 0.969 | 0.940 | 0.982 | 0.965 | 0.943 | 0.988 |
| BACH | 0.977 | 0.988 | 0.998 | 1.000 | 0.996 | 0.959 | 0.998 |
| CCRCC-TCGA-HEL | 0.993 | 0.997 | 0.994 | 0.996 | 0.994 | 0.992 | 0.997 |
| ESCA | 0.886 | 0.894 | 0.899 | 0.903 | 0.902 | 0.889 | 0.902 |
| PanCancer-TIL | 0.946 | 0.975 | 0.965 | 0.977 | 0.971 | 0.949 | 0.978 |
| PanCancer-TCGA | 0.975 | 0.999 | 0.992 | 0.997 | 0.991 | 0.978 | 0.999 |
| UniToPath | 0.830 | 0.838 | 0.836 | 0.840 | 0.865 | 0.823 | 0.844 |
| chaoyang | 0.930 | 0.952 | 0.950 | 0.952 | 0.942 | 0.941 | 0.956 |
| WSSS4LUAD | 0.989 | 0.997 | 0.997 | 0.997 | 0.995 | 0.994 | 0.998 |
| BreakHis | 0.986 | 0.997 | 0.995 | 0.998 | 0.991 | 0.989 | 0.998 |
| PathVQA | 0.574 | 0.560 | 0.603 | 0.613 | 0.628 | 0.594 | 0.613 |
| Report Generation | 0.123 | 0.269 | 0.134 | 0.235 | 0.149 | 0.137 | 0.254 |
| Image Retrieval | 0.892 | 0.938 | 0.883 | 0.958 | 0.943 | 0.874 | 0.967 |
| Average Metric | 0.798 | 0.849 | 0.830 | 0.844 | 0.832 | 0.815 | 0.859 |
| Average Ranking | 6.55 | 2.97 | 4.26 | 2.96 | 4.14 | 5.76 | 1.36 |
| p-value | 3.6e-12 | 1.7e-4 | 1.4e-7 | 3.0e-7 | 8.2e-6 | 3.6e-12 | - |

### v3

@@ 4.4 ROI Classification
For patch-level tissue classification tasks, we evaluate the transfer performance and representation ability of different FMs using a linear probe, inspired by the approach employed in DINOv2. [48, 84]. Initially, we extract features from the images using the pretrained FMs. Subsequently, we employ a linear layer for performing classification. To optimize the model, we utilize AdamW [85] with an initial learning rate of 5e-4 and weight decay of 1e-5. Additionally, we incorporate a cosine annealing scheduler to update the learning rate during training [86]. In order to obtain the best model, we set the maximum number of epochs to 3000 and implemented early stopping with patience of 100 epochs. For ensuring fair comparison, we maintain a consistent batch size of 256 across all methods.
To evaluate the performance of patch-level tissue classification, we consider the impact of class imbalance in the dataset and assess the metrics of balanced accuracy, weighted F1 score, and AUC. These metrics provide comprehensive insights into the classification performance, accounting for both accuracy and the ability to handle imbalanced class distributions. Specifically, we compare the FMs across 16 tasks. For all experiments in this section, we estimate the model performance using non-parametric bootstrapping with 1,000 bootstrap replicates. We employ Torchmetrics [87] for bootstrapping sampling and obtain the mean and standard deviation of the metrics. The experimental results are presented in Extended Data Table 25 to Extended Data Table 36. Furthermore, we report the average performance of the patch-level tissue classification results across 12 tasks in Table 24, demonstrating the superior performance of GPFM.
@@
The experimental results are presented in Extended Data Table 27.
@@
TIL classification (2 classes).We use PanCancer-TIL dataset [94, 95] for tumor infiltrating lymphocyte (TIL) classification. It includes 304,097 images with a size of 100100 pixels at 0.5 micrometers per pixel.
@@
Additionally, we also adopted the data from Center-3 to conduct external validation.
The TIL-negative samples (8,361 ROIs) were obtained from healthy lymph nodes of pan-cancer type, and TIL-positive samples (10,131 ROIs) were obtained from the marked cancerous areas on the lymph nodes with metastasis.
@@
Additionally, we adopted 9,214 normal ROIs, and 11,854 adenoma ROIs from Center-3 for external validation.
The experimental results are presented in Extended Data Table 35.
GasHisDB for Gastric Tissue Classification (2 classes). The dataset consists of a total of 13,124 160160 abnormal images, and 20,160 normal images. For training and evaluation, we resize all patches to 224224 pixels and label-stratified the train-test with a ratio of 8:2 (26,627: 6,657 ROIs). Additionally, we adopted the 709 normal tissues and 1,828 abnormal tissues from Center-3 to perform external validation. Results can be found in Extended Data Table 36.
4.5 Pathological Tissue Retrieval
In the linear probe evaluation tasks, we extract semantically-rich features using different FMs and then construct a task-specific classifier. These features are not only applicable for supervised learning but also prove to be valuable for image-to-image retrieval. The primary goal of this application is to retrieve images that share the same class label as a given query image, thereby facilitating efficient image retrieval. The CRC-100K dataset comprises 100,000 non-overlapping 224224 patches extracted from 86 human cancer tissue slides stained with H&E for training purposes. Additionally, it includes 7,180 images with 224224 pixels extracted from 50 patients diagnosed with colorectal adenocarcinoma for testing. The dataset consists of multiple classes, including Adipose (ADI, 11,745 ROIs), Background (BACK, 11,413 ROIs), Debris (DEB, 11,851 ROIs), Lymphocytes (LYM, 12,191 ROIs), Mucus (MUC, 9,931 ROIs), Smooth muscle (MUS, 14,128 ROIs), Normal colon mucosa (NORM, 9,504 ROIs), Cancer-associated stroma (STR, 10,867 ROIs), and Colorectal adenocarcinoma epithelium (TUM, 15,550 ROIs). For training and evaluation, we utilize the official train-test split, with 100,000 samples for training and 7,180 samples for testing.
To initiate the pathological tissue image retrieval process, we begin by embedding all images using pretrained FMs. Next, each image in the test set is treated as a query and compared against the images in the training set. To ensure that all features have a comparable impact on the computation of similarity, we independently normalize each feature component to the range [0, 1] [101]. This normalization process involves calculating the mean and variance of the training set features, which are then used to normalize both the training and testing features.
@@ 4.6 Pathology Visual Question Answering
The objective of this subsection is to evaluate the performance of our proposed pathology FM in the context of Visual Question Answering (VQA) tasks. To this end, we utilized the PathVQA dataset [102] and the WSI-VQA dataset [51] as benchmark datasets for our experiments. These datasets provide a comprehensive framework for assessing the model’s ability to comprehend and reason about both patch-level and WSI-level visual pathology information, enabling accurate responses to queries related to observed pathological features.
Patch-level VQA on PathVQA dataset.
To evaluate the effectiveness of FMs in pathology VQA, we utilize the PathVQA dataset [102], which is the largest and most widely used dataset in the pathology domain for VQA tasks. The dataset consists of 32,799 image-question-answer triplets, divided into three subsets: a training set (50%) containing 16,400 triplets used for model training, a validation set (30%) comprising 9,840 triplets for hyperparameter tuning and overfitting prevention, and a test set (20%) including 6,560 triplets for final model performance evaluation.
To ensure a rigorous and comparative analysis, we adopt the Multi-modal Unified Medical Captioning (MUMC) method [103], which currently represents the state-of-the-art approach on the PathVQA dataset. The MUMC method has exhibited superior performance in leveraging the synergies between visual and textual information for medical image understanding tasks.
The VQA model architecture consists of four main components: the image encoder, text encoder, multimodal encoder, and answering decoder. The image encoder is responsible for capturing domain-specific visual features. We employ various pathology FMs as the image encoder. During the fine-tuning process, the weights of the image encoder are kept frozen to preserve the integrity of the pre-trained visual representations and focus on learning task-specific multimodal interactions.
@@
WSI-level VQA on WSI-VQA dataset.
The dataset comprises 977 WSIs and 8,671 question-answering pairs, which are divided into three subsets: training, validation, and test.
Specifically, the training subset consists of 804 WSIs and 7,139 pairs, while the validation subset includes 87 WSIs and 798 pairs. The test subset contains 86 WSIs and 735 pairs. In the close-ended portion of the test subset, the correct answers are distributed as follows: 151 for option A, 107 for B, 86 for C, and 46 for D.
For the WSI-VQA dataset, we adhere to the implementation framework proposed by Chen et al. [51], with modifications limited to replacing the visual features.
@@ 4.7 Pathology Report Generation
The task of pathology report generation is inspired by existing works on Chest X-ray and other medical report generation [105, 106, 107]. In this task, the report generation model takes a WSI as input and generates the corresponding pathology report. Specifically, the input WSI is first processed by FMs to extract an initial representation. This representation is then fed into the encoder-decoder architecture of report generation models to produce the decoded pathology report. During this process, the visual encoder further processes the initial representations of WSIs through specific designs [106, 107, 52] to obtain the optimal WSI features for the report decoding stage. The text decoder of the model then utilizes these features for report generation. A good initial representation of WSI could significantly facilitate both the visual encoding and textual decoding stages. Consequently, the quality of the generated report is directly influenced by the representations provided by the FMs. In this task, we adopt the HistGen model [52] for WSI report generation and set the learning rate to 1e-4, and weight decay to 0.8 per epoch. The model is trained for 40 epochs with batch size 1 using features extracted from different FMs.
To evaluate the report generation performance of FMs, we utilize natural language generation metrics including BLEU [108], METEOR [109], and ROUGE-L [110], in which BLEU is further split into BLEU-1, BLEU-2, BLEU-3, and BLEU-4 for evaluation of different granularity. These metrics provide a robust framework for evaluating machine-generated text, each bringing unique strengths to assess different aspects of text quality. This task is conducted on the TCGA WSI-Report dataset proposed in [52] containing 7,690 WSIs and the paired diagnosis reports in total, and the PatchGastricADC dataset [55] which includes 991 pairs of histological descriptions and WSIs of stomach adenocarcinoma endoscopic biopsy specimens. A 7:1:2 train-validation-test split is employed and the experimental results are reported in Extended Data Table 40 and 42.
To assess the robustness of each FM in report generation, we conducted a stratified analysis of the TCGA WSI-Report dataset based on cancer types, focusing on major organ cancers including breast, lung, and kidney. The stratified evaluation results are presented in Table 41. Additionally, we collaborated with an experienced pathologist to perform a rigorous human evaluation of the reports generated by different models. The evaluation employed a four-tier scoring system (illustrated in Fig. 11d), and the scoring distribution and average score of each FM are summarized in Table 43.
@@ 4.8 Computing Software and Hardware
In this project, we utilized PyTorch [111] (version 2.1.2 with CUDA 12.1) for both pretraining and evaluating downstream tasks. To pretrain the GPFM model, we incorporated established FMs, namely UNI, Phikon, and CONCH, as additional teachers. It is worth noting that access to UNI and CONCH requires a prior application submission. The GPFM model was pretrained using the FullyShardedDataParallel (FSDP) technique on 28 80GB NVIDIA H800 GPU nodes. All other data processing and evaluation for downstream tasks were carried out on a server equipped with 8 NVIDIA RTX 3090 GPUs. To assess the model’s performance, we employed Torchmetrics [87] and Scikit-learn [112] for metric evaluation. For WSI processing, we relied on openslide-python (version 1.2.0) [57] and the CLAM [58] codebase. Pathology VQA evaluation was conducted using the MUMC [103] codebase. Furthermore, for histology report generation, we utilized the HistGen [52] codebase. Please refer to Extended Data Table 51 for a comprehensive list of the aforementioned models and libraries utilized in this study.
@@ 5 Data availability
This study incorporates a total of 56 datasets. Out of these, 33 datasets are utilized for pretraining, and a subset of them is also employed for evaluation purposes. The remaining 23 datasets are specifically dedicated to downstream task evaluation. For detailed information on the public data used in this project, refer to the Extended Data Table 52. For the data from Center-1 to Center-5, these datasets are not publicly available due to patient privacy obligations, institutional review board requirements, and data use agreements. However, researchers interested in accessing de-identified data may submit a reasonable request directly to the corresponding authors, subject to obtaining the necessary ethical approvals and complying with institutional policies. The splits of the dataset can be found in our GitHub repository.
@@ 7 Ethics declarations
This project has been reviewed and approved by the Human and Artefacts Research Ethics Committee (HAREC) of Hong Kong University of Science and Technology. The protocol number is HREP-2024-0212.
@@ 8 Author contributions
J.M. conceived the study and designed the experiments. J.M., Z.G., and F.Z. collected the data for self-supervised learning and downstream task evaluation. J.M. performed model pretraining and conducted patch-level tissue classification tasks. Y.L. and X.J. participated in discussions regarding the design of the self-supervised learning framework and were responsible for reproducing the foundation models. J.M., F.Z., and Y.C. evaluated the weakly-supervised WSI classification task. Z.G. performed survival analysis and report generation tasks. Y.W. and Y.X. conducted pathological image retrieval and curated the data of WSI-report pairs. Z.Z. performed pathological image visual question answering. C.J. assisted in result analysis and the creation of visualized attention maps. J.M. and Z.G. prepared the manuscript with input from all co-authors. R.C.K.C, A.H., and L.L. provided medical guidance. L.L., J.L., C.Z., D.L. provided and preprocessed data for some downstream tasks. S.Z. and F.Y. provided preprocessed data for external validation. P.F., J.W. offered insightful suggestions for the experimental design and thoughtfully directing the research trajectory. K.T.C reviewed and refined the draft. H.C. supervised the research.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62202403), Hong Kong Innovation and Technology Commission (Project No. PRP/034/22FX and No. ITCPD/17-9), Research Grants Council of the Hong Kong Special Administrative Region, China (No. R6003-22) and HKUST Frontier Technology Research for Joint Institutes with Industry (No. OKT24EG01). We thank the support of HKUST SuperPod for providing the GPU platform for foundation model pretraining.